### v1

# Efficient Input-Constrained Impulsive Optimal Control of Linear Systems with Application to Spacecraft Relative Motion

Abstract
This work presents a novel algorithm for impulsive optimal control of linear time-varying systems with the inclusion of input magnitude constraints. Impulsive optimal control problems, where the optimal input solution is a sum of delta functions, are typically formulated as an optimization over a normed function space subject to integral equality constraints and can be efficiently solved for linear time-varying systems in their dual formulation. In this dual setting, the problem takes the form of a semi-infinite program which is readily solvable in online scenarios for constructing maneuver plans. This work augments the approach with the inclusion of magnitude constraints on the input over time windows of interest, which is shown to preserve the impulsive nature of the optimal solution and enable efficient solution procedures via semi-infinite programming. The resulting algorithm is demonstrated on the highly relevant problem of relative motion control of spacecraft in Low Earth Orbit (LEO) and compared to several other proposed solutions from the literature.
I Introduction
The problem of minimizing the norm of a control input function space of a linear time-varying system with boundary conditions has received significant attention due to its practical application to many systems of interest. Such optimal control problems arise in many fields, including astrodynamics [1], epidemiology [2], robotics [3], and economics [4]. Despite its widespread applicability, impulsive optimal control remains challenging for many applications of interest, which may involve complex state and input constraints and require efficient and reliable solution procedures. For example, in the context of spacecraft relative motion control, for which computational resources are limited, safety is essential, and algorithmic reliability is crucial, few existing approaches are adequate.
Approaches to solving impulsive optimal control problems are numerous, and include direct methods, such as sequential convex programming [5], indirect methods, leveraging moment problems [6], reachable set theory [7], or primer-vector theory [1], and other techniques, like particle swarm optimization [8] and analytical methods [9]. In direct methods, the dynamics and cost are discretized over a time grid of interest and solved as finite-dimensional optimization problem using tools like convex optimization [10]. These techniques have a tendency to introduce a large number of optimization variables and constraints and can produce suboptimal policies in the case of coarse gridding. Remedies to this issue include having impulse times as optimization variables [5], but this results in a non-linear program that must be solved through techniques like sequential convex programming [11], which may not necessarily provide convergence guarantees [12]. Despite this, these techniques are highly desirable due to their generality and ability to handle input and state constraints. In contrast to direct methods, analytical methods directly solve optimality conditions to derive analytical expressions for control signals but suffer from limited applicability as the solutions can be complicated and specific to a given system. However, they are highly reliable and efficient, and have seen success in the context of spacecraft control [9]. Finally, other techniques like particle swarm optimization [8] and genetic algorithms [13] have also been studied but are typically not well equipped for efficiency and generalizability.
Indirect methods, in comparison, involve formulating optimality conditions on the continuous-time problem, and include moment-based formulations, reachable set theory, and primer-vector theory approaches, which provide different perspectives to the problem and yet all arrive at equivalent optimality conditions. Primer-vector theory, originally coined by Lawden [14] in 1963, derives first order necessary conditions of optimality which can be derived from Pontryagin’s Maximum Principle. The moment problem dates back to the 1960s, when Neustadt [6] showed that the optimal control problem over normed integrable function spaces with integral equality constraints can be embedded in a conjugate space and that the corresponding function space is a sum of delta functions. In the 1970s, Gilbert [7] approached the problem from the reachable set theory perspective, formulating necessary conditions for optimality in terms of spacecraft reachable sets to derive efficient algorithms for fuel-optimal spacecraft rendezvous.
Early algorithmic frameworks for leveraging the optimality conditions derived from these indirect methods are proposed in Gilbert [7], which employed an early numerical algorithm based on quadratic programming [15] to efficiently solve impulsive rendezvous problems. Other approaches involve simplified techniques to solve the necessary conditions to suboptimality [16] or applying mixed iterative schemes to solving the polynomial necessary and sufficient conditions of optimality [17]. In 2013, Arzelier [18] connected the moment problem theorized by Neustadt [6] to a semi-infinite convex optimization scheme [19] to produce an efficient, convergent, and reliable strategy to solving the linearized impulsive optimal control problem. The discretization strategy proposed in this work has served as the foundation for many impulsive optimal control studies for spacecraft relative motion control, such as [20], which connected the technique to reachable set theory and extended the technique to time varying cost functions, and [21], which proposed an extended algorithm to handle state constraints and input constraints. This semi-infinite program (SIP) approach has seen great success in the context of spacecraft rendezvous and relative motion control due to its efficiency and reliability, with heritage on CubeSat missions like VISORS [22].
Despite the unilateral utility of these theoretical and algorithmic frameworks, the applicability of indirect methods to constrained problems remains limited. With respect to spacecraft rendezvous, a major constraint of interest involves bounding the maximum impulse that a spacecraft can execute at a given time. All satellites, such as those of the VISORS mission, have propulsion systems that naturally constrain the amount of $\Delta V$ that can be applied to the system at a given time, necessitating the use of maneuver splitting algorithms or other heuristic approaches [22] which lead to suboptimal maneuver plans and generally increased computational burden. More broadly, it is often practical to distribute $\Delta V$ expenditure for spacecraft relative motion irrespective of propulsion constraints to reduce the aggressiveness of maneuver plans. In the case of rendezvous over 10s of kilometers, naively applying a scheme without the consideration of impulse magnitude can result in maneuver plans that are unsafe, unrealizable, or otherwise undesirable.
To address this theoretical and algorithmic gap, this paper introduces impulse magnitude constraints by applying inequality constraints over the integral of input normed function spaces. It is shown that the resulting formulation preserves the impulsive quality of the solution, and can be easily extended to semi-infinite programming techniques to produce an efficient and reliable solution procedure. The algorithm is demonstrated for several optimal control problems of interest in spacecraft relative motion control and compared with existing approaches.
II Problem Background
This work considers linear, time-varying systems of the form
$\dot{\bm{x}}(t)=A(t)\bm{x}(t)+B(t)\bm{u}(t),$ (1)
where $\bm{x}:[t_{0},t_{f}]\to\mathbb{R}^{n_{x}},t\to\bm{x}(t)$ is the state trajectory represents the system state, $\bm{u}:[t_{0},t_{f}]\to\mathbb{R}^{n_{u}},t\to\bm{u}(t)$ is the input signal, $A:[t_{0},t_{f}]\to\mathbb{R}^{n_{x}\times n_{x}}$ is the piecewise-continuous plant matrix, and $B:[t_{0},t_{f}]\to\mathbb{R}^{n_{x}\times n_{u}}$ is the input matrix. Boundary conditions are enforced at fixed initial and terminal times such that
$\bm{x}(t_{0})=\bm{x}_{0},\ \bm{x}(t_{f})=\bm{x}_{f},$ (2)
where $\bm{x}_{0}\in\mathbb{R}^{n_{x}}$ and $\bm{x}_{f}\in\mathbb{R}^{n_{x}}$ are fixed initial and terminal constraints. The set of input functionals that satisfy the system dynamics and boundary conditions can be expressed as an equality constraint in terms of the state transition matrix via
$\bm{x}_{f}=\Phi(t_{f},t_{0})\bm{x}_{0}+\int_{t_{0}}^{t_{f}}\Phi(t_{f},\tau)B(\tau)\bm{u}(\tau)d\tau.$ (3)
Impulsive optimal control problems typically seek to minimize an accumulated norm of the control input, i.e.,
$J(\bm{u}(\cdot))=\int_{t_{0}}^{t_{f}}f(\bm{u}(\tau))d\tau,$ (4)
where $f:\mathbb{R}^{n_{u}}\to\mathbb{R}^{+}$ is a norm that satisfies positive definiteness ($f(\bm{u})\geq 0\ \forall\bm{u}\in\mathbb{R}^{n_{u}}$, homogeneity ($f(\alpha\bm{u})=\alpha f(\bm{u})\ \forall\alpha\in\mathbb{R}^{+}$), and the triangle inequality ($f(\bm{u}_{1})+f(\bm{u}_{2})\geq f(\bm{u}_{1}+\bm{u}_{2})\ \forall\bm{u}_{1},\bm{u}_{2}\in\mathbb{R}^{n_{u}}$).
Problem 1 (Basic Impulsive Control)
Find $\bm{u}(\cdot)$ to minimize $J(\bm{u}(\cdot))$ (Equation 4) such that the dynamics (Equation 3) are satisfied.
Problem 1 is a moment problem of the same form given in Neustadt [6] whose optimal solution is a sum of delta functions, i.e. the input is an atomic measure. In its primal form, this problem is difficult to solve due to its infinite dimensional nature. The dual form of the problem can be derived from the Lagrangian, which takes the form
$\mathcal{L}(\bm{u}(\cdot),$ $\bm{\lambda}_{f})=\bm{\lambda_{f}}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})$ (5)
$+\int_{t_{0}}^{t_{f}}f(\bm{u}(\tau))-\bm{\lambda}_{f}^{\intercal}\Phi(t_{f},\tau)B(\tau)u(\tau)\text{d}\tau,$
where $\bm{\lambda}_{f}$ is a Lagrangian multiplier associated with the equality constraint in Equation 3 and has a dual function given by
$g(\bm{\lambda}_{f})=\inf_{\bm{u}(\cdot)}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f})$ (6)
[…]
Here, $f^{\circ}(\cdot):\mathbb{R}^{n_{u}}\to\mathbb{R}^{+}$ is the dual norm of $f$, i.e. $f^{\circ}(\bm{y})=\sup_{f(\bm{x})\leq 1}\bm{x}^{\intercal}\bm{y}$. The dual form of the problem is arrived to by the program formed by maximizing the dual function, i.e.
Problem 2 (Basic Dual Problem)
$\mathcal{D}=\max_{\bm{\lambda}_{f}}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})$ (7)
s.t. $f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1\quad\forall t\in[t_{0},t_{f}]$
Problem 7 consists of maximizing an affine function subject to a semi-infinite constraint over time over the norm of the Lagrangian multiplier. As such, it is convex.
In contrast to the primal problem, which consists of minimization of a function space, and is therefore infinite dimensional, the dual problem is an optimization problem over a finite number of variables subject to an infinite constraint space. The dual form of this problem is in fact easier to solve in this form, as Arzelier [18] showed, through leveraging standard discretization methods for semi-infinite programs.
The moment problem is connected to both primer-vector theory and reachable set theory approaches through its dual form. Indeed, $\bm{\lambda}_{v}(t)\triangleq B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$ is in fact the classic primer-vector [14] which satisfies $\dot{\bm{\lambda}}(t)=-A^{\intercal}(t)\bm{\lambda}(t),\bm{\lambda}_{v}(t)=B^{\intercal}(t)\bm{\lambda}(t)$. In this sense, the Lagrangian multiplier $\bm{\lambda}_{f}$ introduced in Equation 5 is simply the costate at the terminal time and the semi-infinite constraint in Problem 7 is equivalent to that derived in Prussing [1] ( $\lambda_{v}^{\intercal}\lambda_{v}\leq 1$ in the case of the $L_{2}$ norm). The same comparison can be made for the more geometric interpretation of reachable set theory [7], which is formulated in terms of support functions of reachable sets. The dual norm of the costate, $f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$, is simply the support function of the sublevel set of the unit cost at time $t$, and $\max_{t\in[t_{0},t_{f}]}f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$ is nothing but the support function of the pseudostate reachable set under unit cost [20].
Problem 3 (Basic Maneuver Reconstruction)
Let $\bm{\lambda}_{f}^{*}$ be the optimal solution to Problem 7 and $t^{*}=\{t\in[t_{0},t_{f}]\mid f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})=1\}$. The set $t^{*}$ represents the optimal impulse times to the system. Neustadt [6] showed that the number of impulse times is at most $n_{x}$, and that selecting $n_{x}$ components of $t^{*}$ recovers the optimal semi-infinite variables. From these times, the optimal control can be constructed as $\bm{u}(t)=\sum_{t_{k}\in t^{*}}\bm{v}_{k}\delta(t-t_{k})$ where $\{\bm{v}_{k}\}_{k=1}^{n}$ satisfies
$\bm{x}_{f}=\Phi(t_{f},t_{0})\bm{x}_{0}+\sum_{k=1}^{n}\Phi(t_{f},t_{k})B(t_{k})\bm{v}_{k}$ (8)
III Problem Formulation
The novel contribution of this paper is the inclusion of input magnitude constraints on the control input. The consideration of input magnitude constraints is non-trivial, since simply enforcing that the norm of the control input be below a given value naturally breaks the impulsivity of the solution. Moreover, enforcing that any given impulse be below some impulse bound does not necessarily preserve the discretizability of the semi-infinite program.
The following modification to Problem 1 is proposed:
Problem 4 (Magnitude Constrained Problem)
Find $\bm{u}(\cdot)$ that solves the primal optimization problem $\mathcal{P}=$
$\min_{\bm{u}(\cdot)}$ $\sum_{k=1}^{N}\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t$ (9)
s.t. $\bm{x}_{f}=\Phi(t_{f},t_{0})\bm{x}_{0}+\sum_{k=1}^{N}\int_{T_{k}}\Phi(t_{f},t)B(t)\bm{u}(t)\text{d}t,$
$\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t\leq\Delta V_{max,k}\quad\forall k\in\mathbb{Z}_{1}^{N},$
where $f_{k}(\cdot)$ are norms, $\Delta V_{max,k}$ are fixed impulse magnitude limits, and $T_{k}$ represent ordered control input time windows that satisfy
$\bigcup_{k=1}^{N}T_{k}\subseteq[t_{0},t_{f}],\ T_{i}\cap T_{j}=\emptyset\ \forall i\neq j.$ (10)
III-A Dual Formulation
To show that Problem 10 is amenable to being solved through semi-infinite programming, the dual form can be derived in a similar manner to Problem 1. The Lagrangian of Equation 9 is
$\begin{gathered}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})=\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})\\ +\sum_{k=1}^{N}\int_{T_{k}}(1+\sigma_{k})f_{k}(\bm{u}(t))-\bm{\lambda}_{f}^{\intercal}\Phi(t_{f},t)B(t)\bm{u}(t)\text{d}t\\ -\Delta V_{max,k}\sigma_{k},\end{gathered}$ (11)
and its dual function is given by
$\begin{gathered}g(\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})=\inf_{\bm{u}(\cdot)}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})\\ =\begin{cases}\bm{\lambda_{f}}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k},\\ \text{if}\ f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t)f,t)\bm{\lambda}_{f})\leq 1\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N},\\ -\infty,\text{otherwise}\end{cases}\end{gathered}$ (12)
From this, the dual form is arrived to with
Problem 5 (Magnitude Constrained Dual Problem)
Find $\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N}$ to solve $\mathcal{D}=$
$\max_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k}$ (13)
s.t. $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1+\sigma_{k}$
$\forall t\in T_{k},\ \forall k\in\mathbb{Z}_{1}^{N}$
This problem is indeed familiar to the unconstrained form of Problem 7. The problem consists still of maximizing an affine function subject to a semi-infinite constraint over the dual norm of the costate. Moreover, the problem has been reduced from an optimization over an infinite function space in the primal to an optimization over $n_{x}+N$ variables with $N$ semi-infinite constraints in the dual. This dual form permits the dual norm violation of $1$, but penalizes this violation in the cost weighted by the associated maximum impulse magnitude.
Several theorems are now presented to show that the dual form retains several desirable properties.
Theorem III.1
Problem 13 is strongly dual to the primal Problem 10, i.e. $\mathcal{D}=\mathcal{P}$.
Proof:
Evidently, Problem 13 satisfies Slater’s condition for semi-infinite programs (see Definition 2.1 from [23]) since there is a feasible interior point, i.e. $\exists\bm{\lambda}_{f},\bm{\sigma}>0,\text{s.t.}f_{k}^{\circ}(B^{\intercal}(t)\Phi(t_{f},t)\bm{\lambda}_{f})<1+\sigma_{k}\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N}$. From Theorem 2.3 of [23], this implies that strong duality holds for the semi-infinite program. ∎
This result is important for establishing feasibility of the primal problem, as it implies that when the dual problem is unbounded above ($\mathcal{D}=\infty$), then the primal problem is infeasible, and no input exists to steer the dynamical system between the boundary conditions subject to the input magnitude constraints.
Theorem III.2
The semi-infinite dual Problem 13 is reducible [23], i.e. there exists a finite subset of times $\{t\}\triangleq\{t_{1},\dots,t_{M}\}\subset\bigcup_{k=1}^{N}T_{k}$ such that
$\max_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k}$ (14)
s.t. $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1+\sigma_{k}$
$\forall t\in T_{k}\cap\{t\},\ \forall k\in\mathbb{Z}_{1}^{N}$
provides an equivalent solution to Problem 13.
Proof:
Since the problem exhibits strong duality, the problem is reducible so long as an optimal solution exists (the primal is feasible) according to Theorem 3.1 of [23]. ∎
This property is important in both the theoretical and algorithmic sense. With respect to the former, it shows that the optimal solution still takes the form of a sum of delta functions, meaning the solution structure is preserved under the formulation of Problem 10. With respect to the latter, discretization techniques can be applied to the semi-infinite dual problem to efficiently determine the solution of the primal problem.
III-B Input Reconstruction
The technique used to recover the optimal input to the unconstrained problem from the dual formulation can be similarly applied to its constrained version. If $\bm{\lambda}_{f}^{*},\bm{\sigma}^{*}$ are the optimal solution variables to Problem 13, then the optimal impulse times for a given window are recovered from the set
$\{t\}^{*}_{k}$ $=\{t\in T_{k}\mid f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})=1+\sigma_{k}^{*}\}$ (15)
$=\{t_{k1},t_{k2},\dots,t_{k,m_{k}}\}.$
and the corresponding input is given by
$\bm{u}^{*}(t)=\sum_{k=1}^{N}\sum_{j=1}^{m_{k}}\bm{v}_{kj}\delta(t-t_{kj}).$ (16)
Furthermore, from the duality formulation, it must hold that
$\sigma_{k}\left(\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t-\Delta V_{max,k}\right)=0\ \forall k\in\mathbb{Z}_{1}^{N},$ (17)
which can be leveraged to show the interesting property that
$\sigma_{k}>0$ $\iff\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=\Delta V_{max,k}$ (18)
$\iff\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})=\Delta V_{max,k},$
meaning that the non-zero elements of $\bm{\sigma}$ correspond to time windows in which the input is saturated.
From the optimal impulse times, the optimal input can be recovered from the problem:
Problem 6 (Constrained Input Reconstruction Problem)
Solve the discretized primal Problem 10 from the optimal impulse times recovered from Problem 13 with $\mathcal{P}_{d}=$
$\min_{\{\bm{v}\}}$ $\sum_{k=1}^{N}\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})$ (19)
s.t. $\footnotesize{\bm{x}_{f}=\Phi(t_{f},t_{0})\bm{x}_{0}+\sum_{k=1}^{N}\sum_{j=1}^{m_{k}}\Phi(t_{f},t_{kj})B(t_{kj})\bm{v}_{kj},}$
$\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})\leq\Delta V_{max,k}\qquad\forall k\in\mathbb{Z}_{1}^{N}.$
IV Algorithmic Approach
Solving Problem 10 in an algorithmically tractable manner can be done through first solving the dual problem to determine the optimal impulse locations and then reconstructing the input through the discretized primal. However, certain nuances must be considered to solve these problems efficiently and robustly.
IV-A Dual Problem Algorithm
Problem 13 in its native form is not readily solvable due to the semi-infinite constraint. Several techniques exist to solve semi-infinite programs (SIPs), but this work and typical applications to spacecraft rendezvous leverage discretization methods, which convert the problem to an iterative finite optimization problem. Moreover, in contrast to Problem 7, which has no possibility of being unbounded above, Problem 13 can be unbounded, corresponding to primal infeasibility. When applying the iterative discretization technique, it is therefore necessary to ensure that premature unboundedness does not produce artificial infeasibility of the problem.
To illustrate the iterative technique to solving the semi-infinite problem, consider the finite form of Problem 13, given by $\mathcal{D}(S)=$
$\max_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k}$ (20)
s.t. $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1+\sigma_{k}$
$\forall t\in S\cap T_{k},\ \forall k\in\mathbb{Z}_{1}^{N}$
$\|\bm{\lambda}_{f}\|_{2}\leq\lambda_{max}$
where $S\triangleq\{s_{1},s_{2},\dots s_{d}\}\subset\bigcup_{k=1}^{N}T_{k}$ is a finite set. Since Problem 13 is reducible according to Theorem III.2, then it holds that the finite version of the problem can solve the semi-infinite problem. To leverage this property, discretization techniques gradually add values to the set $S$ until $\mathcal{D}(S)\to\mathcal{D}$. This can be done recursively at each solution iteration, given the optimal solution of the previous iteration $\bm{\lambda}_{f}^{(j)},\bm{\sigma}^{(j)},S^{(j)}$ with
$S^{(j+1)}\leftarrow S^{(j)}\cup\bigcup_{k=1}^{N}\arg\max_{t\in T_{k}}f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)}).$ (21)
The discrete set can be initialized simply with $S^{(1)}=\bigcup_{k=1}^{N}\partial T_{k}$ or with the optimal times associated with a two-impulse burn as is done in [18].
The finite optimization problem in Equation 20 also includes a norm constraint $\|\bm{\lambda}_{f}\|_{2}\leq\lambda_{max}$ to protect against artificial infeasibility in the case that the discretized problem is unbounded. Upon algorithmic convergence, the activeness of this constraint determines whether the semi-infinite problem is unbounded, in which case the primal is infeasible.
These algorithmic modifications retain several important features of discretization methods. Namely, it is evident that $\mathcal{D}\leq\mathcal{D}(S^{(j+1)})\leq\mathcal{D}(S^{(j)})$ since $S^{(j)}\subseteq S^{(j+1)}$. Moreover, given a tolerance $\epsilon$, an important result is as follows, which is extended from [18].
Lemma 1
Suppose Equation 20 is solved such that $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})-\sigma_{k}^{*}\leq 1+\epsilon\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N}$. Then it follows that $\mathcal{D}(S^{(j)})\leq(1+\epsilon)\mathcal{D}$.
Proof:
Let $\bar{\bm{\lambda}}_{f}=\bm{\lambda}_{f}^{*}/(1+\epsilon)$ and $\bar{\sigma}_{k}=\sigma_{k}^{*}/(1+\epsilon)$. Evidently, $(\bar{\bm{\lambda}}_{f},\bar{\bm{\sigma}})$ is a feasible point to Problem 13 that satisfies $\mathcal{D}(S^{(j)})/(1+\epsilon)=\bar{\bm{\lambda}}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\bar{\sigma}_{k}\leq\mathcal{D}\implies\mathcal{D}(S^{(j)})\leq(1+\epsilon)\mathcal{D}$. ∎
Lemma 1 can be exploited in an algorithmic framework to ensure that the dual problem is solved to within a specified suboptimality tolerance. Algorithm 10 summarizes the procedure for solving the dual problem.
0: LTV Dynamics $\{\Phi(\tau_{f},\tau_{0}),A(\tau),B(\tau)\}$, Time Intervals $\{T_{k}\}_{k=1}^{N}$, Norms $\{f_{k}(\cdot)\}_{k=1}^{N}$ Boundary Conditions $\{\bm{x}_{0},\bm{x}_{f}\}$, Convergence Tolerance $\epsilon$
0: Optimal Cost $\mathcal{D}(S^{(j)})$, Optimal Values $\bm{\lambda}_{f},\bm{\sigma}$
1: Discretize $\{T_{k}\}_{k=1}^{N}$
[…]
3:while $\exists k\in\mathbb{Z}_{1}^{N},t\in T_{k}\ \text{s.t.}\ f^{\circ}_{k}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})-\sigma_{k}^{*}>1+\epsilon$ do
4:  $\bm{\lambda}_{f}^{(j)},\bm{\sigma}^{(j)},\mathcal{D}(S^{(j)})\leftarrowSolve$ Optimization Problem 20 with $S^{(j)}$
5:  $S^{(j+1)}\leftarrow S^{(j)}\cup\bigcup_{k=1}^{N}\arg\max_{t\in T_{k}}f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)})$
6:  $j\leftarrow j+1$
7:end while
8:if $\|\bm{\lambda}_{f}^{(j)}\|_{2}=\lambda_{max}$ then
9:  return Problem Infeasible
10:end if
Algorithm 1 Dual Semi-Infinite Program Algorithm.
IV-B Input Reconstruction
Reconstructing the input signal from the dual solution requires extracting a discrete set of times where the semi-infinite constraints are active and solving the discretized form of the primal problem. These aspects are summarized in Algorithm 2.
0: LTV Dynamics $\{\Phi(\tau_{f},\tau_{0}),A(\tau),B(\tau)\}$, Time Intervals $\{T_{k}\}_{k=1}^{N}$, Norms $\{f_{k}(\cdot)\}_{k=1}^{N}$ Boundary Conditions $\{\bm{x}_{0},\bm{x}_{f}\}$, Dual Solution $\{\bm{\lambda}_{f}^{*},\bm{\sigma}^{*}\}$
0: Optimal Cost $\mathcal{P}_{d}$, Input Signal $\bm{u}(t)$, State $\bm{x}(t)$
[…]
Algorithm 2 Input Reconstruction Algorithm.
V Performance and Validation
The problem formulation and corresponding algorithmic approach is validated on the highly relevant problem of impulsive control for proximity operations of satellites.
V-A System Definition
Relative motion of satellites in orbits subject to single body gravitational force can be modeled as linear time-varying systems under a variety of additional perturbations, such as drag or spherical harmonics, through a relative orbital element state representation [24, 25], defined by
$\delta\textit{{\oe }}=\begin{bmatrix}\delta a&\delta\lambda&\delta e_{x}&\delta e_{y}&\delta i_{x}&\delta i_{y}\end{bmatrix}^{\intercal}$. (22)
These states are functions of the classical Keplerian orbital elements of circular two-body motion of the target and chaser spacecraft. Notably, they represent relative motion in terms of integration constants and convert the system to its Jordan canonical form. The linearized equations of motion under this state representation are given by
[…]
Here, $n\triangleq\sqrt{\frac{\mu}{a^{3}}}$ is the orbital mean motion defined based on the planetary gravitational constant ($\mu=398,600.4418\text{km}^{3}/\text{s}^{2}$ for Earth) and the semi-major axis of the orbit ($a=[6538,8378]$ km for LEO). Additionally, $u_{0}$ is the initial mean argument of latitude, which is set to 0 for all ensuing examples. Under these equations of motion, the state transition matrix is simply
$\Phi(t_{f},t_{0})=A(t_{f}-t_{0})+I_{6\times 6}$ (24)
Finally, relative orbital elements can be linearly transformed to relative position and relative velocities in the radial-transverse-normal frame of the target spacecraft using the similarity transformation given by $\delta\bm{x}(t)\triangleq[x,y,z,v_{x},v_{y},v_{z}]^{\intercal}=R(t)\delta\textit{{\oe }}(t)$ where
$R(t)=\begin{bmatrix}1&0&-\text{c}(q)&-\text{s}(q)&0&0\\ 0&1&2\text{s}(q)&-2\text{c}(q)&0&0\\ 0&0&0&0&\text{s}(q)&-\text{c}(q)\\ 0&0&\text{s}(q)&-\text{c}(q)&0&0\\ -\frac{3n}{2}&0&2\text{c}(q)&2\text{s}(q)&0&0\\ 0&0&0&0&\text{c}(q)&\text{s}(q)\end{bmatrix}$ (25)
and $q=nt+u_{0}$.
V-B Planar Reconfiguration Demonstration
The first example considers a planar rendezvous example consisting of reducing along track separation from a distance of 100 km to 10 km. The initial and final conditions are given in Table I. The transfer occurs over four orbits, with eight equally spaced time windows enforced over the duration of the transfer with a buffer of $1/10$th of an orbit between each window. Each time window is discretized over 1000 time points in the dual optimization scheme to provide sufficient resolution. The input is scaled such that a unitary impulse corresponds to $1.13$ [m/s] of $\Delta V$ expenditure. Moreover, the norm applied is the $L_{2}$ norm, such that $f_{k}(\cdot)=\|\cdot\|_{2}\ \forall k$. Finally, the orbital mean motion is $n=.00113$.
TABLE I: Initial Conditions for Planar Reconfiguration (km)
$a\delta a$ $a\delta\lambda$ $a\delta e_{x}$ $a\delta e_{y}$ $a\delta i_{x}$ $a\delta i_{y}$
$\delta\textit{{\oe }}_{0}$ 10 100 0 10 0 0
$\delta\textit{{\oe }}_{f}$ 0 10 0 0 0 0
(a) Relative Position
(b) Throttle
(c) Dual Norm
(d) $\bm{\lambda}_{v}(t)$
Figure 1: Planar Reconfiguration Demonstration.
(a) $\delta a-\delta\lambda$ Trajectory
(b) $\delta e_{x}-\delta e_{y}$ Trajectory
Figure 2: Planar Reconfiguration Relative Orbital Elements
Figure 1 displays the trajectory corresponding to the computed planar reconfiguration control signal and Figure 2 shows the trajectory from the perspective of the relative orbital elements. The computed dual optimal solution for the problem is $\bm{\lambda}_{f}^{*}=[0.7380,0.0891,-0.2359,-0.7236,0,0]^{\intercal},\bm{\sigma}^{*}=[5.4503,3.0092,3.7714,1.3608,2.0926,0,0.4142,1.5183]^{\intercal}$. The total $\Delta V$ expenditure is $\Delta V_{max}\sum_{k=1}^{N}\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=8.817$ [m/s]. Figure 1a shows the relative positions on the radial-transverse plane with impulses occurring at the blue dots. Figure 1b displays the throttle profile of the computed input signal in blue and time windows shown in red. Evidently, the time windows where the input signal is saturated, i.e. $\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=1$, correspond to when $\sigma_{k}$ is non-zero. Figure 1c displays the dual norm associated with the semi-infinite constraint and the times at which the dual norm is maximized for each time window. The dual norm exceeds the value of 1 at all but one of the time windows, corresponding exactly to the time windows at which the input signal is saturated. Finally, Figure 1d plots the trajectory formed by $B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$, which equates to the primer vector $\bm{\lambda}_{v}(t)$ [14], and corresponding impulse locations in blue. Additionally, the unit circle is plotted, corresponding to the sublevel set of the semi-infinite constraint at $\sigma=0$.
To validate and compare the performance of the proposed algorithm, the problem is also solved under a direct optimization framework whereby the primal problem is discretized with the inputs at each discrete node as solution variables. Under this approach, the number of solution variables drastically increases compared to the dual solution framework, with the total number of solution variables amounting to $n_{u}NM$, with $n_{u}$ being the input dimension, $N$ the number of time windows, and $M$ the discretization count per time window. In contrast, the dual SIP involves $n_{x}+N$ solution variables, with cone count being on the order of $\sim 2N$-$3N$, depending on the problem structure. The poor scalability of direct primal optimization manifests as longer solver times, with this issue being especially apparent for long duration transfers, whereby the discretization induces substantial variable counts. This issue is especially prevalent for real-time optimal control of CubeSats, for which CPU performance and memory are limited.
Under this direct optimization approach, the equivalent problem is solved and produces an equivalent solution. However, the solver clock time is 0.6 seconds, whereas the cumulative solver time for the dual SIP with input reconstruction is 0.1262 seconds. Runtime experiments were performed on an Intel i7‑10750H CPU in MATLAB R2024a with the MOSEK solver [26] and YALMIP modeler [27]. Solver times are determined from extracting the MOSEK solver time from the YALMIP solution. Runtime disparities between the proposed approach and direct primal optimization are likely to be more severe for longer duration transfers, where node counts are large, and less severe for shorter duration transfers.
V-C VISORS Transfer Demonstration
The VISORS mission [22] is a two-nanosatellite mission that aims to perform telescopic measurements of the solar corona. Part of the mission operations involves performing a long duration, impulse constrained transfer from a 200 meter radius standby orbit to a 40 meter science orbit. The challenge of this transfer is confounded by the limitations of the propulsion systems aboard the nanosatellites, which can provide a maximum impulse of just 2 millimeters per second with at least 45 seconds of cool down time between each impulse. The current technique used to realize this transfer involves applying closed-form solvers to a sequence of manually specified waypoints that gradually tightens the relative orbit. The main issues with this approach are that it is inflexible to incorporating dynamical perturbations, it does not provide optimal solutions, and it is not adaptable to differing initial and final conditions.
To demonstrate the strength of the proposed algorithm for this difficult problem, the standby to science scenario from [22] is implemented. Initial and final conditions for this transfer are given in Table II. The transfer duration is 10 orbits and 500 equally spaced time windows are enforced with a buffer of 45 seconds and a maximum impulse expenditure of $2$ millimeters per second per window. The $L_{1}$ norm is applied such that $f_{k}(\cdot)=\|\cdot\|_{1}\ \forall k$. Each time window is discretized over 100 time points. Finally, the orbital mean motion is $n=.00113$.
TABLE II: Initial Conditions for VISORS Reconfiguration (m)
$a\delta a$ $a\delta\lambda$ $a\delta e_{x}$ $a\delta e_{y}$ $a\delta i_{x}$ $a\delta i_{y}$
$\delta\textit{{\oe }}_{0}$ 0 0 0 200 0 200
$\delta\textit{{\oe }}_{f}$ -2.62 45.21 -34.51 4.78 -18.72 2.72
(a) Relative Position
(b) $B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$
Figure 3: VISORS Reconfiguration Trajectory
(a) Throttle Profile
(b) Dual Norm Profile
Figure 4: VISORS Reconfiguration Input Profile
(a) $\delta a-\delta\lambda$
(b) $\delta\bm{e},\delta\bm{i}$
(c) RN Separation
Figure 5: VISORS Reconfiguration Relative Orbital Elements
The resulting trajectory is displayed in Figure 3, the input profile in Figure 4, and the relative orbital element perspective in Figure 5. In Figure 3b, the trajectory of $B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$, equivalent to the primer vector, is plotted along with the unit cube, which corresponds to the unit sublevel set of the dual norm. Times when the primer vector leaves this region correspond to possible impulse events. The dual norm of the primer vector is shown in Figure 4b, along with time windows and impulse times, which occur when the dual norm signal is maximized over the time window. The relative orbital elements reveal an important feature of interest of the computed maneuver sequence, namely that RN separation, plotted in Figure 5c and typically used to evaluate the safe spacecraft separation [28], is maintained above 18 meters for the entire transfer duration, despite the fact that no state constraints are applied in the solution procedure.
The maneuver sequence under the proposed algorithm requires 0.3095 meters per second of total $\Delta V$ expenditure. In contrast, [22] reports 0.4934 meters per second of $\Delta V$ expenditure for the same transfer. Given that this transfer is performed 20 times over the mission span, this amounts to a total $\Delta V$ saving of 3.68 meters per second, which is $\sim$25% of the total $\Delta V$ budget of the CubeSat. The proposed algorithm is additionally capable of incorporating $J_{2}$ and drag perturbations through the linear model formulations in [25], does not require manual tuning of waypoints along the trajectory, and is flexible to changing initial and final conditions. The total solver clock time is 0.2675 seconds on a modern computer, which is reasonably achievable in an online manner under the lower computational resources of the VISORS CubeSats.
VI Conclusion
This paper demonstrated an extension to a class of moment problems that included constraints on the integral of the norm of the input signal over time windows of interest. It was shown that the resulting optimal solution remains impulsive and that semi-infinite programming can be applied to global optimality and global convergence. This algorithmic framework was applied to challenging spacecraft relative motion control problems, where it was used to derive maneuver sequences that outperformed established techniques in terms of optimality.
Though this paper has focused on applications to linear time-varying systems, the authors believe that the theoretical framework outlined has significant promise for extending indirect trajectory optimization techniques for impulsively controlled nonlinear systems. Future work could also include the incorporation of state constraints and extensions to impulsive covariance control.

### v2

@@
 This work augments the approach with the inclusion of magnitude constraints on the input over time windows of interest, which is shown to preserve the impulsive nature of the optimal solution and enable efficient solution procedures via semi-infinite programming.
-The resulting algorithm is demonstrated on the highly relevant problem of relative motion control of spacecraft in Low Earth Orbit (LEO) and compared to several other proposed solutions from the literature.
+The resulting algorithm is demonstrated on the highly relevant problem of relative motion control of spacecraft in Low Earth Orbit (LEO).
 I Introduction
 The problem of minimizing the norm of a control input function space of a linear time-varying system with boundary conditions has received significant attention due to its practical application to many systems of interest.
-Such optimal control problems arise in many fields, including astrodynamics [1], epidemiology [2], robotics [3], and economics [4].
+Such optimal control problems arise in many fields and are central to the study of spacecraft optimal control and trajectory optimization.
 Despite its widespread applicability, impulsive optimal control remains challenging for many applications of interest, which may involve complex state and input constraints and require efficient and reliable solution procedures.
 For example, in the context of spacecraft relative motion control, for which computational resources are limited, safety is essential, and algorithmic reliability is crucial, few existing approaches are adequate.
-Approaches to solving impulsive optimal control problems are numerous, and include direct methods, such as sequential convex programming [5], indirect methods, leveraging moment problems [6], reachable set theory [7], or primer-vector theory [1], and other techniques, like particle swarm optimization [8] and analytical methods [9].
-In direct methods, the dynamics and cost are discretized over a time grid of interest and solved as finite-dimensional optimization problem using tools like convex optimization [10].
-These techniques have a tendency to introduce a large number of optimization variables and constraints and can produce suboptimal policies in the case of coarse gridding.
-Remedies to this issue include having impulse times as optimization variables [5], but this results in a non-linear program that must be solved through techniques like sequential convex programming [11], which may not necessarily provide convergence guarantees [12].
-Despite this, these techniques are highly desirable due to their generality and ability to handle input and state constraints.
-In contrast to direct methods, analytical methods directly solve optimality conditions to derive analytical expressions for control signals but suffer from limited applicability as the solutions can be complicated and specific to a given system.
-However, they are highly reliable and efficient, and have seen success in the context of spacecraft control [9].
-Finally, other techniques like particle swarm optimization [8] and genetic algorithms [13] have also been studied but are typically not well equipped for efficiency and generalizability.
-Indirect methods, in comparison, involve formulating optimality conditions on the continuous-time problem, and include moment-based formulations, reachable set theory, and primer-vector theory approaches, which provide different perspectives to the problem and yet all arrive at equivalent optimality conditions.
-Primer-vector theory, originally coined by Lawden [14] in 1963, derives first order necessary conditions of optimality which can be derived from Pontryagin’s Maximum Principle.
-The moment problem dates back to the 1960s, when Neustadt [6] showed that the optimal control problem over normed integrable function spaces with integral equality constraints can be embedded in a conjugate space and that the corresponding function space is a sum of delta functions.
-In the 1970s, Gilbert [7] approached the problem from the reachable set theory perspective, formulating necessary conditions for optimality in terms of spacecraft reachable sets to derive efficient algorithms for fuel-optimal spacecraft rendezvous.
+Approaches to solving impulsive optimal control problems are numerous, and include direct methods, such as sequential convex programming [18], indirect methods, leveraging moment problems [19], reachable set theory [10], or primer-vector theory [20], and other techniques, such as analytical methods [7].
+In direct methods, the dynamics and cost are discretized over a time grid of interest and solved as a finite-dimensional optimization problem.
+Despite widespread utility, these techniques do not explicitly exploit solution sparsity and therefore tend to introduce many optimization variables or can produce suboptimal solutions in the case of coarse gridding.
+Indirect methods, in comparison, involve formulating optimality conditions on the continuous-time problem, and include moment-based formulations [19], reachable set theory [10], and primer-vector theory approaches [16], which provide different perspectives to the problem and yet all arrive at equivalent optimality conditions.
 Early algorithmic frameworks for leveraging the optimality conditions derived from these indirect methods are proposed in Gilbert [10], which employed an early numerical algorithm based on quadratic programming [5] to efficiently solve impulsive rendezvous problems.
 Other approaches involve simplified techniques to solve the necessary conditions to suboptimality [6] or applying mixed iterative schemes to solving the polynomial necessary and sufficient conditions of optimality [3].
-In 2013, Arzelier [18] connected the moment problem theorized by Neustadt [6] to a semi-infinite convex optimization scheme [19] to produce an efficient, convergent, and reliable strategy to solving the linearized impulsive optimal control problem.
-The discretization strategy proposed in this work has served as the foundation for many impulsive optimal control studies for spacecraft relative motion control, such as [20], which connected the technique to reachable set theory and extended the technique to time varying cost functions, and [21], which proposed an extended algorithm to handle state constraints and input constraints.
-This semi-infinite program (SIP) approach has seen great success in the context of spacecraft rendezvous and relative motion control due to its efficiency and reliability, with heritage on CubeSat missions like VISORS [22].
-Despite the unilateral utility of these theoretical and algorithmic frameworks, the applicability of indirect methods to constrained problems remains limited.
+In 2013, Arzelier [4] connected the moment problem theorized by Neustadt [19] to a semi-infinite programming convex optimization scheme leveraging exchange methods [21] to produce an efficient, convergent, and reliable strategy to solving the linearized impulsive optimal control problem.
+This semi-infinite programming approach has served as the foundation for subsequent impulsive optimal control studies [14, 13] and has heritage on CubeSat missions such as VISORS [11].
+Despite the practicality of these theoretical and algorithmic frameworks, the applicability of indirect methods to constrained problems remains limited.
 With respect to spacecraft rendezvous, a major constraint of interest involves bounding the maximum impulse that a spacecraft can execute at a given time.
-All satellites, such as those of the VISORS mission, have propulsion systems that naturally constrain the amount of $\Delta V$ that can be applied to the system at a given time, necessitating the use of maneuver splitting algorithms or other heuristic approaches [22] which lead to suboptimal maneuver plans and generally increased computational burden.
+All satellites, such as those of the VISORS mission, have propulsion systems that naturally constrain the amount of $\Delta V$ that can be applied at a given time, necessitating the use of maneuver splitting algorithms or other heuristic approaches [11] which lead to suboptimal maneuver plans or increased computational burden.
 More broadly, it is often practical to distribute $\Delta V$ expenditure for spacecraft relative motion irrespective of propulsion constraints to reduce the aggressiveness of maneuver plans.
-In the case of rendezvous over 10s of kilometers, naively applying a scheme without the consideration of impulse magnitude can result in maneuver plans that are unsafe, unrealizable, or otherwise undesirable.
-To address this theoretical and algorithmic gap, this paper introduces impulse magnitude constraints by applying inequality constraints over the integral of input normed function spaces.
-It is shown that the resulting formulation preserves the impulsive quality of the solution, and can be easily extended to semi-infinite programming techniques to produce an efficient and reliable solution procedure.
-The algorithm is demonstrated for several optimal control problems of interest in spacecraft relative motion control and compared with existing approaches.
+To address this theoretical and algorithmic gap, this paper introduces impulse magnitude constraints by applying inequality constraints to the integral of the input norm over specified time windows.
+It is shown that the resulting formulation preserves the impulsive quality of the solution and can be easily extended to semi-infinite programming techniques to yield an efficient and reliable solution procedure.
+The algorithm is demonstrated for several optimal control problems of interest in spacecraft relative motion control.
+In contrast to other formulations of input-constrained optimal control [1, 23], which are ubiquitous in aerospace literature, the formulation presented here is unique in that it preserves solution impulsivity, allowing for both improved algorithmic procedures and more desirable maneuver plans.
 II Problem Background
@@
 where $\bm{x}_{0}\in\mathbb{R}^{n_{x}}$ and $\bm{x}_{f}\in\mathbb{R}^{n_{x}}$ are fixed initial and terminal constraints.
-The set of input functionals that satisfy the system dynamics and boundary conditions can be expressed as an equality constraint in terms of the state transition matrix via
+The set of input functions that satisfy the system dynamics and boundary conditions can be expressed as an equality constraint in terms of the state transition matrix via
 $\bm{x}_{f}=\Phi(t_{f},t_{0})\bm{x}_{0}+\int_{t_{0}}^{t_{f}}\Phi(t_{f},\tau)B(\tau)\bm{u}(\tau)d\tau.$ (3)
@@
 Find $\bm{u}(\cdot)$ to minimize $J(\bm{u}(\cdot))$ (Equation 4) such that the dynamics (Equation 3) are satisfied.
-Problem 1 is a moment problem of the same form given in Neustadt [6] whose optimal solution is a sum of delta functions, i.e.
+Problem 1 is a moment problem of the same form given in Neustadt [19], whose optimal solution is a sum of delta functions, i.e.
 the input is an atomic measure.
 In its primal form, this problem is difficult to solve due to its infinite dimensional nature.
-The dual form of the problem can be derived from the Lagrangian, which takes the form
-$\mathcal{L}(\bm{u}(\cdot),$ $\bm{\lambda}_{f})=\bm{\lambda_{f}}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})$ (5)
-$+\int_{t_{0}}^{t_{f}}f(\bm{u}(\tau))-\bm{\lambda}_{f}^{\intercal}\Phi(t_{f},\tau)B(\tau)u(\tau)\text{d}\tau,$
-where $\bm{\lambda}_{f}$ is a Lagrangian multiplier associated with the equality constraint in Equation 3 and has a dual function given by
-$g(\bm{\lambda}_{f})=\inf_{\bm{u}(\cdot)}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f})$ (6)
-[…]
-Here, $f^{\circ}(\cdot):\mathbb{R}^{n_{u}}\to\mathbb{R}^{+}$ is the dual norm of $f$, i.e.
+However, in the dual setting, the problem is conveniently reformulated as the following convex semi-infinite program
+Problem 2 (Basic Dual Problem)
+$\sup_{\bm{\lambda}_{f}}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})$ (5)
+s.t. $f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1\quad\forall t\in[t_{0},t_{f}]$
+where the optimization variable $\bm{\lambda}_{f}$ is a Lagrange multiplier and $f^{\circ}(\cdot):\mathbb{R}^{n_{u}}\to\mathbb{R}^{+}$ is the dual norm of $f$, i.e.
 $f^{\circ}(\bm{y})=\sup_{f(\bm{x})\leq 1}\bm{x}^{\intercal}\bm{y}$.
-The dual form of the problem is arrived to by the program formed by maximizing the dual function, i.e.
-Problem 2 (Basic Dual Problem)
-$\mathcal{D}=\max_{\bm{\lambda}_{f}}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})$ (7)
-s.t. $f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1\quad\forall t\in[t_{0},t_{f}]$
-Problem 7 consists of maximizing an affine function subject to a semi-infinite constraint over time over the norm of the Lagrangian multiplier. As such, it is convex.
-In contrast to the primal problem, which consists of minimization of a function space, and is therefore infinite dimensional, the dual problem is an optimization problem over a finite number of variables subject to an infinite constraint space.
-The dual form of this problem is in fact easier to solve in this form, as Arzelier [18] showed, through leveraging standard discretization methods for semi-infinite programs.
+For many relevant norms, e.g., $l_{p}$ norms, the dual norm is analytically available ($\|\cdot\|_{p}^{\circ}=\|\cdot\|_{q},1/p+1/q=1$, see also [14]).
+In contrast to the infinite dimensional primal problem, the dual problem is an optimization problem over a finite number of variables subject to an infinite constraint space.
+The dual form of this problem is in fact easier to solve in this form, as Arzelier [4] leveraged, through exchange methods for semi-infinite programs [21, 12].
 The moment problem is connected to both primer-vector theory and reachable set theory approaches through its dual form.
-Indeed, $\bm{\lambda}_{v}(t)\triangleq B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$ is in fact the classic primer-vector [14] which satisfies $\dot{\bm{\lambda}}(t)=-A^{\intercal}(t)\bm{\lambda}(t),\bm{\lambda}_{v}(t)=B^{\intercal}(t)\bm{\lambda}(t)$.
-In this sense, the Lagrangian multiplier $\bm{\lambda}_{f}$ introduced in Equation 5 is simply the costate at the terminal time and the semi-infinite constraint in Problem 7 is equivalent to that derived in Prussing [1] ( $\lambda_{v}^{\intercal}\lambda_{v}\leq 1$ in the case of the $L_{2}$ norm).
-The same comparison can be made for the more geometric interpretation of reachable set theory [7], which is formulated in terms of support functions of reachable sets.
-The dual norm of the costate, $f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$, is simply the support function of the sublevel set of the unit cost at time $t$, and $\max_{t\in[t_{0},t_{f}]}f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$ is nothing but the support function of the pseudostate reachable set under unit cost [20].
+Indeed, $\bm{\lambda}_{v}(t)\triangleq B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$ is in fact the classic primer-vector [16], with the Lagrange multiplier $\bm{\lambda}_{f}$ being the costate at the terminal time, and the semi-infinite constraint in Problem 5 is equivalent to that derived in Prussing [20].
+From the perspective of reachable set theory [10], the dual norm of the costate, $f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$, is simply the support function of the sublevel set of the unit cost at time $t$, and $\max_{t\in[t_{0},t_{f}]}f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})$ is nothing but the support function of the pseudostate reachable set under unit cost [14].
 Problem 3 (Basic Maneuver Reconstruction)
-Let $\bm{\lambda}_{f}^{*}$ be the optimal solution to Problem 7 and $t^{*}=\{t\in[t_{0},t_{f}]\mid f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})=1\}$.
+Let $\bm{\lambda}_{f}^{*}$ be the optimal solution to Problem 5 and $t^{*}=\{t\in[t_{0},t_{f}]\mid f^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})=1\}$.
 The set $t^{*}$ represents the optimal impulse times to the system.
@@
 The novel contribution of this paper is the inclusion of input magnitude constraints on the control input.
-The consideration of input magnitude constraints is non-trivial, since simply enforcing that the norm of the control input be below a given value naturally breaks the impulsivity of the solution.
-Moreover, enforcing that any given impulse be below some impulse bound does not necessarily preserve the discretizability of the semi-infinite program.
+The consideration of input magnitude constraints is nontrivial, since bounded control magnitudes, as are typically enforced in aerospace literature for continuous thrust formulations [1, 23], prevents the optimal solution from concentrating mass at discrete time points.
+As will be shown, the formulation presented here maintains impulsivity of the input profile, which provides both algorithmic benefits and a desirable input structure for spacecraft which are often modeled as impulsively actuated.
 The following modification to Problem 1 is proposed:
 Problem 4 (Magnitude Constrained Problem)
-Find $\bm{u}(\cdot)$ that solves the primal optimization problem $\mathcal{P}=$
+Find $\bm{u}(\cdot)$ that solves the primal optimization problem $\mathcal{P}(\{T_{k}\}_{k=1}^{N})=$
 $\min_{\bm{u}(\cdot)}$ $\sum_{k=1}^{N}\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t$ (7)
 s.t. $\bm{x}_{f}=\Phi(t_{f},t_{0})\bm{x}_{0}+\sum_{k=1}^{N}\int_{T_{k}}\Phi(t_{f},t)B(t)\bm{u}(t)\text{d}t,$
-$\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t\leq\Delta V_{max,k}\quad\forall k\in\mathbb{Z}_{1}^{N},$
-where $f_{k}(\cdot)$ are norms, $\Delta V_{max,k}$ are fixed impulse magnitude limits, and $T_{k}$ represent ordered control input time windows that satisfy
-$\bigcup_{k=1}^{N}T_{k}\subseteq[t_{0},t_{f}],\ T_{i}\cap T_{j}=\emptyset\ \forall i\neq j.$ (10)
+$\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t\leq\Delta V_{k}\quad\forall k\in\mathbb{Z}_{1}^{N},$
+where $f_{k}(\cdot)$ are norms, $\Delta V_{k}$ are fixed impulse magnitude limits, and $T_{k}$ represent ordered control input time windows that satisfy $\bigcup_{k=1}^{N}T_{k}\subseteq[t_{0},t_{f}],\ T_{i}\cap T_{j}=\emptyset\ \forall i\neq j.$
 III-A Dual Formulation
-To show that Problem 10 is amenable to being solved through semi-infinite programming, the dual form can be derived in a similar manner to Problem 1. The Lagrangian of Equation 9 is
-$\begin{gathered}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})=\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})\\ +\sum_{k=1}^{N}\int_{T_{k}}(1+\sigma_{k})f_{k}(\bm{u}(t))-\bm{\lambda}_{f}^{\intercal}\Phi(t_{f},t)B(t)\bm{u}(t)\text{d}t\\ -\Delta V_{max,k}\sigma_{k},\end{gathered}$ (11)
-and its dual function is given by
-$\begin{gathered}g(\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})=\inf_{\bm{u}(\cdot)}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})\\ =\begin{cases}\bm{\lambda_{f}}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k},\\ \text{if}\ f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t)f,t)\bm{\lambda}_{f})\leq 1\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N},\\ -\infty,\text{otherwise}\end{cases}\end{gathered}$ (12)
+To show that Problem 4 is amenable to being solved through semi-infinite programming, the dual form can be derived in a similar manner to that used for Problem 1. The Lagrangian of Equation 7 is
+$\begin{gathered}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\bm{\sigma})=\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\sigma_{k}\\ +\sum_{k=1}^{N}\int_{T_{k}}(1+\sigma_{k})f_{k}(\bm{u}(t))-\bm{\lambda}_{f}^{\intercal}\Phi(t_{f},t)B(t)\bm{u}(t)\text{d}t,\end{gathered}$ (8)
+where $\bm{\lambda}_{f}$ and $\bm{\sigma}=(\sigma_{k})_{k=1}^{N}$ are Lagrange multipliers associated with the equality and inequality constraints of the primal in Problem 4 respectively.
+The associated dual function is given by
+$\begin{gathered}g(\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})=\inf_{\bm{u}(\cdot)}\mathcal{L}(\bm{u}(\cdot),\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N})\\ =\begin{cases}\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\sigma_{k},\\ \text{if}\ f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1+\sigma_{k}\ \\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N};-\infty,\text{otherwise}\end{cases}\end{gathered}$ (9)
 From this, the dual form is arrived to with
 Problem 5 (Magnitude Constrained Dual Problem)
-Find $\bm{\lambda}_{f},\{\sigma_{k}\}_{k=1}^{N}$ to solve $\mathcal{D}=$
-$\max_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k}$ (13)
+Find $\bm{\lambda}_{f},\bm{\sigma}$ such that $\mathcal{D}(\{T_{k}\}_{k=1}^{N})=$
+$\sup_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\sigma_{k}$ (10)
 s.t. $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1+\sigma_{k}$
 $\forall t\in T_{k},\ \forall k\in\mathbb{Z}_{1}^{N}$
-This problem is indeed familiar to the unconstrained form of Problem 7.
-The problem consists still of maximizing an affine function subject to a semi-infinite constraint over the dual norm of the costate.
-Moreover, the problem has been reduced from an optimization over an infinite function space in the primal to an optimization over $n_{x}+N$ variables with $N$ semi-infinite constraints in the dual.
-This dual form permits the dual norm violation of $1$, but penalizes this violation in the cost weighted by the associated maximum impulse magnitude.
-Several theorems are now presented to show that the dual form retains several desirable properties.
-Theorem III.1
-Problem 13 is strongly dual to the primal Problem 10, i.e. $\mathcal{D}=\mathcal{P}$.
+Similar to Problem 5, the problem has been reduced from an optimization over an infinite function space in the primal to a convex semi-infinite optimization problem over $n_{x}+N$ variables.
+Lemma 1
+Problem 10 is strongly dual to the primal Problem 4, i.e. $\mathcal{D}=\mathcal{P}$.
 Proof:
-Evidently, Problem 13 satisfies Slater’s condition for semi-infinite programs (see Definition 2.1 from [23]) since there is a feasible interior point, i.e.
+Evidently, Problem 10 satisfies Slater’s condition for semi-infinite programs (see Definition 2.1 from [22]) since there is a feasible interior point, i.e.
 $\exists\bm{\lambda}_{f},\bm{\sigma}>0,\text{s.t.}f_{k}^{\circ}(B^{\intercal}(t)\Phi(t_{f},t)\bm{\lambda}_{f})<1+\sigma_{k}\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N}$.
@@
 ∎
-This result is important for establishing feasibility of the primal problem, as it implies that when the dual problem is unbounded above ($\mathcal{D}=\infty$), then the primal problem is infeasible, and no input exists to steer the dynamical system between the boundary conditions subject to the input magnitude constraints.
-Theorem III.2
-The semi-infinite dual Problem 13 is reducible [23], i.e. there exists a finite subset of times $\{t\}\triangleq\{t_{1},\dots,t_{M}\}\subset\bigcup_{k=1}^{N}T_{k}$ such that
-$\max_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k}$ (14)
-s.t. $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1+\sigma_{k}$
-$\forall t\in T_{k}\cap\{t\},\ \forall k\in\mathbb{Z}_{1}^{N}$
-provides an equivalent solution to Problem 13.
+Lemma 2
+The semi-infinite dual Problem 10 is reducible [22], i.e., there exists a finite subset of times $\{t\}\triangleq\{t_{1},\dots,t_{M}\}\subset\bigcup_{k=1}^{N}T_{k}$ such that $\mathcal{D}(\{t\})=\mathcal{D}(\{T_{k}\}_{k=1}^{N})$.
 Proof:
 Since the problem exhibits strong duality, the problem is reducible so long as an optimal solution exists (the primal is feasible) according to Theorem 3.1 of [22]. ∎
-This property is important in both the theoretical and algorithmic sense.
-With respect to the former, it shows that the optimal solution still takes the form of a sum of delta functions, meaning the solution structure is preserved under the formulation of Problem 10.
-With respect to the latter, discretization techniques can be applied to the semi-infinite dual problem to efficiently determine the solution of the primal problem.
+Lemmas 1 and 2 extend several important properties to the proposed constrained problem.
+Namely, unboundedness of the dual problem implies infeasibility of the primal, the optimal solution retains impulsivity (i.e., $\bm{u}(t)=0$ a.e.), and exchange methods [21] are valid for solving Problem 10.
 III-B Input Reconstruction
 The technique used to recover the optimal input to the unconstrained problem from the dual formulation can be similarly applied to its constrained version.
-If $\bm{\lambda}_{f}^{*},\bm{\sigma}^{*}$ are the optimal solution variables to Problem 13, then the optimal impulse times for a given window are recovered from the set
+If $\bm{\lambda}_{f}^{*},\bm{\sigma}^{*}$ are the optimal solution variables to Problem 10, then the optimal impulse times for a given window are recovered from the set
 $\{t\}^{*}_{k}$ $=\{t\in T_{k}\mid f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})=1+\sigma_{k}^{*}\}$ (11)
@@
 $\bm{u}^{*}(t)=\sum_{k=1}^{N}\sum_{j=1}^{m_{k}}\bm{v}_{kj}\delta(t-t_{kj}).$ (12)
-Furthermore, from the duality formulation, it must hold that
-$\sigma_{k}\left(\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t-\Delta V_{max,k}\right)=0\ \forall k\in\mathbb{Z}_{1}^{N},$ (17)
+Furthermore, from the complementary slackness condition, it must hold that
+$\sigma_{k}\left(\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t-\Delta V_{k}\right)=0\ \forall k\in\mathbb{Z}_{1}^{N},$ (13)
 which can be leveraged to show the interesting property that
-$\sigma_{k}>0$ $\iff\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=\Delta V_{max,k}$ (18)
-$\iff\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})=\Delta V_{max,k},$
+$\sigma_{k}>0\implies\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})=\Delta V_{k}$ (14)
 meaning that the non-zero elements of $\bm{\sigma}$ correspond to time windows in which the input is saturated.
-From the optimal impulse times, the optimal input can be recovered from the problem:
+From the optimal impulse times, the optimal input can be recovered from the tractable finite LP/SOCP:
 Problem 6 (Constrained Input Reconstruction Problem)
-Solve the discretized primal Problem 10 from the optimal impulse times recovered from Problem 13 with $\mathcal{P}_{d}=$
-$\min_{\{\bm{v}\}}$ $\sum_{k=1}^{N}\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})$ (19)
-s.t. $\footnotesize{\bm{x}_{f}=\Phi(t_{f},t_{0})\bm{x}_{0}+\sum_{k=1}^{N}\sum_{j=1}^{m_{k}}\Phi(t_{f},t_{kj})B(t_{kj})\bm{v}_{kj},}$
-$\sum_{j=1}^{m_{k}}f_{k}(\bm{v}_{kj})\leq\Delta V_{max,k}\qquad\forall k\in\mathbb{Z}_{1}^{N}.$
+Solve the discretized primal Problem 4 from the optimal impulse times recovered from Problem 10 with $\mathcal{P}_{d}=\mathcal{P}(\{\{t\}_{k}^{*}\}_{k=1}^{N})$.
+Theorem III.1
+A feasible optimal solution to Problem 4 consists of at most $n_{x}+N$ impulses.
+Proof:
+Since strong duality holds and the dual form is a convex semi-infinite program of $n_{x}+N$ variables, Theorem 3.2 of [22] states that there exists a discretization $S$ which satisfies $|S|\leq n_{x}+N$ such that $\mathcal{D}(S)=\mathcal{D}$, leading to at most $n_{x}+N$ points or atoms in time at which $\bm{u}(t)$ may be nonzero.
+∎
+Remark 1
+The number of impulses can be further reduced when considering the addition of the $\bm{\sigma}\geq 0$ constraints, which reduces the active support of the semi-infinite constraint, leading to at most $n_{x}+N_{a}$ points in time at which the input may be nonzero, where $N_{a}$ is the number of strictly positive $\bm{\sigma}$.
 IV Algorithmic Approach
-Solving Problem 10 in an algorithmically tractable manner can be done through first solving the dual problem to determine the optimal impulse locations and then reconstructing the input through the discretized primal.
-However, certain nuances must be considered to solve these problems efficiently and robustly.
+Solving Problem 4 in an algorithmically tractable manner can be done through first solving the dual problem using exchange methods to determine the optimal impulse locations and then reconstructing the input through the discretized primal.
 IV-A Dual Problem Algorithm
-Problem 13 in its native form is not readily solvable due to the semi-infinite constraint.
-Several techniques exist to solve semi-infinite programs (SIPs), but this work and typical applications to spacecraft rendezvous leverage discretization methods, which convert the problem to an iterative finite optimization problem.
-Moreover, in contrast to Problem 7, which has no possibility of being unbounded above, Problem 13 can be unbounded, corresponding to primal infeasibility.
-When applying the iterative discretization technique, it is therefore necessary to ensure that premature unboundedness does not produce artificial infeasibility of the problem.
-To illustrate the iterative technique to solving the semi-infinite problem, consider the finite form of Problem 13, given by $\mathcal{D}(S)=$
-$\max_{\bm{\lambda}_{f},\bm{\sigma}\geq 0}$ $\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\sigma_{k}$ (20)
-s.t. $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f})\leq 1+\sigma_{k}$
-$\forall t\in S\cap T_{k},\ \forall k\in\mathbb{Z}_{1}^{N}$
-$\|\bm{\lambda}_{f}\|_{2}\leq\lambda_{max}$
-where $S\triangleq\{s_{1},s_{2},\dots s_{d}\}\subset\bigcup_{k=1}^{N}T_{k}$ is a finite set.
-Since Problem 13 is reducible according to Theorem III.2, then it holds that the finite version of the problem can solve the semi-infinite problem.
-To leverage this property, discretization techniques gradually add values to the set $S$ until $\mathcal{D}(S)\to\mathcal{D}$.
-This can be done recursively at each solution iteration, given the optimal solution of the previous iteration $\bm{\lambda}_{f}^{(j)},\bm{\sigma}^{(j)},S^{(j)}$ with
-$S^{(j+1)}\leftarrow S^{(j)}\cup\bigcup_{k=1}^{N}\arg\max_{t\in T_{k}}f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)}).$ (21)
-The discrete set can be initialized simply with $S^{(1)}=\bigcup_{k=1}^{N}\partial T_{k}$ or with the optimal times associated with a two-impulse burn as is done in [18].
-The finite optimization problem in Equation 20 also includes a norm constraint $\|\bm{\lambda}_{f}\|_{2}\leq\lambda_{max}$ to protect against artificial infeasibility in the case that the discretized problem is unbounded.
-Upon algorithmic convergence, the activeness of this constraint determines whether the semi-infinite problem is unbounded, in which case the primal is infeasible.
-These algorithmic modifications retain several important features of discretization methods.
+This work leverages an exchange method [21] for solving the semi-infinite program, which consists of iteratively solving a finite version of Problem 10 and adding constraints corresponding to the worst semi-infinite constraint violation.
+To illustrate this iterative technique, define the modified form of Problem 10, given by
+Problem 7
+Solve the discretized form of Problem 10 $\mathcal{D}(S)$, where $S\triangleq\{s_{1},s_{2},\dots s_{d}\}\subset\bigcup_{k=1}^{N}T_{k}$ is a finite set, with the addition of the constraint:
+$\bm{\lambda}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\sigma_{k}\leq\Delta V_{tot},$ (15)
+where $\Delta V_{tot}=\sum_{k=1}^{N}\Delta V_{k}$ is the total amount of expendable fuel imposed.
+Since Problem 10 is reducible according to Lemma 2, then it holds that the finite version of the problem can solve the semi-infinite problem if $S$ is chosen correctly.
+The additional constraint in Problem 7 prevents artificial infeasibility in the case that the discrete form is under constrained in the exchange method process.
+This constraint is natural to apply since $\mathcal{D}\leq\mathcal{P}\leq\Delta V_{tot}$ if a feasible point exists, meaning there is no loss in feasibility when this constraint is imposed.
+To leverage the reducibility property of the dual, exchange methods gradually add values to the set $S$ such that $\mathcal{D}(S^{(j)})\to\mathcal{D}$ [21, 12].
+Several prominent techniques exist in exchange methods for augmenting the discrete set $S$, which effect problem size growth and convergence rate.
+The technique proposed here simply evaluates the semi-infinite constraint over a grid and adds time points corresponding to the largest constraint violations.
+$\begin{gathered}S^{(j+1)}\leftarrow S^{(j)}\cup\operatorname*{arg\,loc\,max}_{\scriptscriptstyle{t\in\cup_{k=1}^{N}T_{k}:f_{k}^{\circ}(\bm{\lambda}_{v}^{(j)}(t))>1+\sigma_{k}^{(j)}}}f_{k}^{\circ}(\bm{\lambda}_{v}^{(j)}(t))\end{gathered}$ (16)
+Here, $\bm{\lambda}_{v}^{(j)}(t)=B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)}$ is the primer vector at iteration $j$ of the algorithm.
+Since $f_{k}^{\circ}(\bm{\lambda}_{v}^{(j)}(t))$ is trivial to compute in this case, this process does not introduce significant computational burden in the algorithmic framework.
+The discrete set can be initialized with simply the initial and final times, i.e., $S^{(1)}=\{t_{0},t_{f}\}$ or with the optimal times associated with a two-impulse burn as is done in [4].
+The selection of the initial times and the addition of new times are hyperparameters of the algorithmic framework which can be adjusted for specific problems, akin to that done in [14].
+These algorithmic modifications retain several important features of exchange methods.
 Namely, it is evident that $\mathcal{D}\leq\mathcal{D}(S^{(j+1)})\leq\mathcal{D}(S^{(j)})$ since $S^{(j)}\subseteq S^{(j+1)}$.
@@
 Lemma 3
-Suppose Equation 20 is solved such that $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})-\sigma_{k}^{*}\leq 1+\epsilon\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N}$.
+Suppose Problem 7 is solved such that $f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})-\sigma_{k}^{*}\leq 1+\epsilon\ \forall t\in T_{k},\forall k\in\mathbb{Z}_{1}^{N}$.
 Then it follows that $\mathcal{D}(S^{(j)})\leq(1+\epsilon)\mathcal{D}$.
@@
 Let $\bar{\bm{\lambda}}_{f}=\bm{\lambda}_{f}^{*}/(1+\epsilon)$ and $\bar{\sigma}_{k}=\sigma_{k}^{*}/(1+\epsilon)$.
-Evidently, $(\bar{\bm{\lambda}}_{f},\bar{\bm{\sigma}})$ is a feasible point to Problem 13 that satisfies $\mathcal{D}(S^{(j)})/(1+\epsilon)=\bar{\bm{\lambda}}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{max,k}\bar{\sigma}_{k}\leq\mathcal{D}\implies\mathcal{D}(S^{(j)})\leq(1+\epsilon)\mathcal{D}$.
+Evidently, $(\bar{\bm{\lambda}}_{f},\bar{\bm{\sigma}})$ is a feasible point to Problem 10 that satisfies $\mathcal{D}(S^{(j)})/(1+\epsilon)=\bar{\bm{\lambda}}_{f}^{\intercal}(\bm{x}_{f}-\Phi(t_{f},t_{0})\bm{x}_{0})-\sum_{k=1}^{N}\Delta V_{k}\bar{\sigma}_{k}\leq\mathcal{D}\implies\mathcal{D}(S^{(j)})\leq(1+\epsilon)\mathcal{D}$.
 ∎
@@
 0: Optimal Cost $\mathcal{D}(S^{(j)})$, Optimal Values $\bm{\lambda}_{f},\bm{\sigma}$
-1: Discretize $\{T_{k}\}_{k=1}^{N}$
-[…]
-3:while $\exists k\in\mathbb{Z}_{1}^{N},t\in T_{k}\ \text{s.t.}\ f^{\circ}_{k}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{*})-\sigma_{k}^{*}>1+\epsilon$ do
-4:  $\bm{\lambda}_{f}^{(j)},\bm{\sigma}^{(j)},\mathcal{D}(S^{(j)})\leftarrowSolve$ Optimization Problem 20 with $S^{(j)}$
-5:  $S^{(j+1)}\leftarrow S^{(j)}\cup\bigcup_{k=1}^{N}\arg\max_{t\in T_{k}}f_{k}^{\circ}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)})$
-6:  $j\leftarrow j+1$
-7:end while
-8:if $\|\bm{\lambda}_{f}^{(j)}\|_{2}=\lambda_{max}$ then
-9:  return Problem Infeasible
-10:end if
+1: Discretize $\{T_{k}\}_{k=1}^{N}$, $S^{(1)}\leftarrow\{t_{0},t_{f}\}$, $j\leftarrow 1$
+2:while $\exists k\in\mathbb{Z}_{1}^{N},t\in T_{k}\ \text{s.t.}\ f^{\circ}_{k}(B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}^{(j)})-\sigma_{k}^{(j)}>1+\epsilon$ do
+3:  $\bm{\lambda}_{f}^{(j)},\bm{\sigma}^{(j)},\mathcal{D}(S^{(j)})\leftarrowSolve$ Problem 7 with $S^{(j)}$
+4:  $S^{(j+1)}\leftarrow S^{(j)}\cup\Delta S$ (Equation 16), $j\leftarrow j+1$
+5:end while
+6:if $\mathcal{D}(S^{(j)})=\Delta V_{tot}$ then
+7:  return Problem Infeasible
+8:end if
 Algorithm 1 Dual Semi-Infinite Program Algorithm.
@@
 Algorithm 2 Input Reconstruction Algorithm.
+Though the above procedure is the most general, it is possible to exploit further algorithmic performance in the case of smooth norm costs and constraints, in which case the direction of the input is known and the ambiguous input magnitudes can be solved through least squares.
+Theorem IV.1
+Solving the infinite program Problem 4, by means of Algorithm 8 followed by Algorithm 2, is globally convergent to the global minimum to within the imposed tolerance $\epsilon$.
+Proof:
+Since Problem 7 is convex and bounded above, each subproblem of Algorithm 8 will converge.
+Moreover, Algorithm 8 will always terminate since $|S^{(j)}|$ is bounded.
+Finally, since $\mathcal{D}(S^{*})\leq\mathcal{D}(1+\epsilon)$ (Lemma 3), $\mathcal{D}=\mathcal{P}$ (Lemma 1), and $\mathcal{D}(S^{*})=\mathcal{P}(S^{*})$ (strong duality of SOCPs/LPs), $\mathcal{P}(S^{*})\leq\mathcal{P}(1+\epsilon)$.
+∎
 V Performance and Validation
-The problem formulation and corresponding algorithmic approach is validated on the highly relevant problem of impulsive control for proximity operations of satellites.
+The problem formulation and corresponding algorithmic approach is validated on the problem of impulsive control for proximity operations of satellites.
 V-A System Definition
-Relative motion of satellites in orbits subject to single body gravitational force can be modeled as linear time-varying systems under a variety of additional perturbations, such as drag or spherical harmonics, through a relative orbital element state representation [24, 25], defined by
+Relative motion of satellites in orbits subject to single body gravitational force can be modeled as linear time-varying systems under a variety of perturbations, such as drag or spherical harmonics, through a relative orbital element state representation [9, 15], defined by
 $\delta\textit{{\oe }}=\begin{bmatrix}\delta a&\delta\lambda&\delta e_{x}&\delta e_{y}&\delta i_{x}&\delta i_{y}\end{bmatrix}^{\intercal}$. (17)
 These states are functions of the classical Keplerian orbital elements of circular two-body motion of the target and chaser spacecraft.
-Notably, they represent relative motion in terms of integration constants and convert the system to its Jordan canonical form.
-The linearized equations of motion under this state representation are given by
-[…]
-Here, $n\triangleq\sqrt{\frac{\mu}{a^{3}}}$ is the orbital mean motion defined based on the planetary gravitational constant ($\mu=398,600.4418\text{km}^{3}/\text{s}^{2}$ for Earth) and the semi-major axis of the orbit ($a=[6538,8378]$ km for LEO).
-Additionally, $u_{0}$ is the initial mean argument of latitude, which is set to 0 for all ensuing examples.
-Under these equations of motion, the state transition matrix is simply
-$\Phi(t_{f},t_{0})=A(t_{f}-t_{0})+I_{6\times 6}$ (24)
-Finally, relative orbital elements can be linearly transformed to relative position and relative velocities in the radial-transverse-normal frame of the target spacecraft using the similarity transformation given by $\delta\bm{x}(t)\triangleq[x,y,z,v_{x},v_{y},v_{z}]^{\intercal}=R(t)\delta\textit{{\oe }}(t)$ where
-$R(t)=\begin{bmatrix}1&0&-\text{c}(q)&-\text{s}(q)&0&0\\ 0&1&2\text{s}(q)&-2\text{c}(q)&0&0\\ 0&0&0&0&\text{s}(q)&-\text{c}(q)\\ 0&0&\text{s}(q)&-\text{c}(q)&0&0\\ -\frac{3n}{2}&0&2\text{c}(q)&2\text{s}(q)&0&0\\ 0&0&0&0&\text{c}(q)&\text{s}(q)\end{bmatrix}$ (25)
-and $q=nt+u_{0}$.
+The linearized equations of motion under this state are of Jordan form and can be found in [15] along with transformation matrices to RTN frame cartesian state.
+The ensuing examples use an orbital mean motion of $n=.00113$, consistent with LEO.
 V-B Planar Reconfiguration Demonstration
-The first example considers a planar rendezvous example consisting of reducing along track separation from a distance of 100 km to 10 km.
-The initial and final conditions are given in Table I.
+The first example considers a planar rendezvous problem consisting of reducing along track separation from a distance of 100 km to 10 km, i.e., $a\delta\textit{{\oe }}_{0}=[10,100,0,10,0,0]^{\intercal}$ [km] to $a\delta\textit{{\oe }}_{f}=[0,10,0,0,0,0]^{\intercal}$ [km].
 The transfer occurs over four orbits, with eight equally spaced time windows enforced over the duration of the transfer with a buffer of $1/10$th of an orbit between each window.
@@
 Moreover, the norm applied is the $l_{2}$ norm, such that $f_{k}(\cdot)=\|\cdot\|_{2}\ \forall k$.
-Finally, the orbital mean motion is $n=.00113$.
-TABLE I: Initial Conditions for Planar Reconfiguration (km)
-$a\delta a$ $a\delta\lambda$ $a\delta e_{x}$ $a\delta e_{y}$ $a\delta i_{x}$ $a\delta i_{y}$
-$\delta\textit{{\oe }}_{0}$ 10 100 0 10 0 0
-$\delta\textit{{\oe }}_{f}$ 0 10 0 0 0 0
-(a) Relative Position
-(b) Throttle
-(c) Dual Norm
-(d) $\bm{\lambda}_{v}(t)$
-Figure 1: Planar Reconfiguration Demonstration.
-(a) $\delta a-\delta\lambda$ Trajectory
-(b) $\delta e_{x}-\delta e_{y}$ Trajectory
+Figure 1: Planar Reconfiguration Trajectory.
 Figure 2: Planar Reconfiguration Relative Orbital Elements
-Figure 1 displays the trajectory corresponding to the computed planar reconfiguration control signal and Figure 2 shows the trajectory from the perspective of the relative orbital elements.
+Figure 1 displays the RT-plane relative position, primer vector trajectory, input impulse profile, and dual norm trajectory.
+Figure 2 shows the trajectory from the perspective of the planar relative orbital elements.
+Blue dots/arrows denote impulse events, green dots denote starting time, and red dots denote terminal time.
 The computed dual optimal solution for the problem is $\bm{\lambda}_{f}^{*}=[0.7380,0.0891,-0.2359,-0.7236,0,0]^{\intercal},\bm{\sigma}^{*}=[5.4503,3.0092,3.7714,1.3608,2.0926,0,0.4142,1.5183]^{\intercal}$.
-The total $\Delta V$ expenditure is $\Delta V_{max}\sum_{k=1}^{N}\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=8.817$ [m/s].
-Figure 1a shows the relative positions on the radial-transverse plane with impulses occurring at the blue dots.
-Figure 1b displays the throttle profile of the computed input signal in blue and time windows shown in red.
-Evidently, the time windows where the input signal is saturated, i.e.
-$\int_{T_{k}}f_{k}(\bm{u}(t))\text{d}t=1$, correspond to when $\sigma_{k}$ is non-zero.
-Figure 1c displays the dual norm associated with the semi-infinite constraint and the times at which the dual norm is maximized for each time window.
-The dual norm exceeds the value of 1 at all but one of the time windows, corresponding exactly to the time windows at which the input signal is saturated.
-Finally, Figure 1d plots the trajectory formed by $B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$, which equates to the primer vector $\bm{\lambda}_{v}(t)$ [14], and corresponding impulse locations in blue.
-Additionally, the unit circle is plotted, corresponding to the sublevel set of the semi-infinite constraint at $\sigma=0$.
+The total $\Delta V$ expenditure is $8.860$ [m/s].
+Notably, the dual norm exceeds the value of 1 at all but one of the time windows, corresponding exactly to the time windows at which the input constraint is active.
 To validate and compare the performance of the proposed algorithm, the problem is also solved under a direct optimization framework whereby the primal problem is discretized with the inputs at each discrete node as solution variables.
-Under this approach, the number of solution variables drastically increases compared to the dual solution framework, with the total number of solution variables amounting to $n_{u}NM$, with $n_{u}$ being the input dimension, $N$ the number of time windows, and $M$ the discretization count per time window.
+Under this approach, the number of solution variables increases drastically compared to the dual solution framework, with the total number of solution variables amounting to $n_{u}NM$, with $n_{u}$ being the input dimension, $N$ the number of time windows, and $M$ the discretization count per time window.
 In contrast, the dual SIP involves $n_{x}+N$ solution variables, with cone count being on the order of $\sim 2N$-$3N$, depending on the problem structure.
 The poor scalability of direct primal optimization manifests as longer solver times, with this issue being especially apparent for long duration transfers, whereby the discretization induces substantial variable counts.
-This issue is especially prevalent for real-time optimal control of CubeSats, for which CPU performance and memory are limited.
 Under this direct optimization approach, the equivalent problem is solved and produces an equivalent solution.
-However, the solver clock time is 0.6 seconds, whereas the cumulative solver time for the dual SIP with input reconstruction is 0.1262 seconds.
-Runtime experiments were performed on an Intel i7‑10750H CPU in MATLAB R2024a with the MOSEK solver [26] and YALMIP modeler [27].
+However, the solver runtime is 0.6 seconds, whereas the cumulative solver time for the dual SIP with input reconstruction is 0.1262 seconds [^1] Runtime experiments were performed on an Intel i7‑10750H CPU in MATLAB R2024a with the MOSEK solver [2] and YALMIP modeler [17].
 Solver times are determined from extracting the MOSEK solver time from the YALMIP solution.
-Runtime disparities between the proposed approach and direct primal optimization are likely to be more severe for longer duration transfers, where node counts are large, and less severe for shorter duration transfers.
-V-C VISORS Transfer Demonstration
-The VISORS mission [22] is a two-nanosatellite mission that aims to perform telescopic measurements of the solar corona.
-Part of the mission operations involves performing a long duration, impulse constrained transfer from a 200 meter radius standby orbit to a 40 meter science orbit.
-The challenge of this transfer is confounded by the limitations of the propulsion systems aboard the nanosatellites, which can provide a maximum impulse of just 2 millimeters per second with at least 45 seconds of cool down time between each impulse.
-The current technique used to realize this transfer involves applying closed-form solvers to a sequence of manually specified waypoints that gradually tightens the relative orbit.
-The main issues with this approach are that it is inflexible to incorporating dynamical perturbations, it does not provide optimal solutions, and it is not adaptable to differing initial and final conditions.
-To demonstrate the strength of the proposed algorithm for this difficult problem, the standby to science scenario from [22] is implemented.
-Initial and final conditions for this transfer are given in Table II.
-The transfer duration is 10 orbits and 500 equally spaced time windows are enforced with a buffer of 45 seconds and a maximum impulse expenditure of $2$ millimeters per second per window.
-The $L_{1}$ norm is applied such that $f_{k}(\cdot)=\|\cdot\|_{1}\ \forall k$.
+.
+V-C VISORS Transfer Demonstration and Comparison
+The VISORS mission [11] is a two-nanosatellite formation-flying mission that aims to perform telescopic measurements of the solar corona.
+Part of the mission operations involves performing a long duration, impulse constrained transfer from a 200 meter radius standby orbit to a 40 meter science orbit, i.e., $a\delta\textit{{\oe }}_{0}=[0,0,0,200,0,200]^{\intercal}$ [m] to $a\delta\textit{{\oe }}_{f}=[-2.62,45.21,-34.51,4.78,-18.72,2.72]^{\intercal}$ [m].
+The challenge of this transfer is compounded by the limitations of the propulsion systems aboard the nanosatellites, which can provide a maximum impulse of just 2 millimeters per second with at least 45 seconds of cool down time between each impulse.
+To demonstrate the strength of the proposed algorithm for this difficult problem, the standby to science scenario from [11] is demonstrated with a transfer duration of 10 orbits and 500 equally spaced time windows are enforced with a buffer of 45 seconds and a maximum impulse expenditure of $2$ millimeters per second per window.
+The $l_{1}$ norm is applied, $f_{k}(\cdot)=\|\cdot\|_{1}\ \forall k$, to be consistent with the fuel metric in [11].
 Each time window is discretized over 100 time points.
-Finally, the orbital mean motion is $n=.00113$.
-TABLE II: Initial Conditions for VISORS Reconfiguration (m)
-$a\delta a$ $a\delta\lambda$ $a\delta e_{x}$ $a\delta e_{y}$ $a\delta i_{x}$ $a\delta i_{y}$
-$\delta\textit{{\oe }}_{0}$ 0 0 0 200 0 200
-$\delta\textit{{\oe }}_{f}$ -2.62 45.21 -34.51 4.78 -18.72 2.72
-(a) Relative Position
-(b) $B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$
-Figure 3: VISORS Reconfiguration Trajectory
-(a) Throttle Profile
-(b) Dual Norm Profile
-Figure 4: VISORS Reconfiguration Input Profile
-(a) $\delta a-\delta\lambda$
-(b) $\delta\bm{e},\delta\bm{i}$
-(c) RN Separation
-Figure 5: VISORS Reconfiguration Relative Orbital Elements
-The resulting trajectory is displayed in Figure 3, the input profile in Figure 4, and the relative orbital element perspective in Figure 5.
-In Figure 3b, the trajectory of $B^{\intercal}(t)\Phi^{\intercal}(t_{f},t)\bm{\lambda}_{f}$, equivalent to the primer vector, is plotted along with the unit cube, which corresponds to the unit sublevel set of the dual norm.
-Times when the primer vector leaves this region correspond to possible impulse events.
-The dual norm of the primer vector is shown in Figure 4b, along with time windows and impulse times, which occur when the dual norm signal is maximized over the time window.
-The relative orbital elements reveal an important feature of interest of the computed maneuver sequence, namely that RN separation, plotted in Figure 5c and typically used to evaluate the safe spacecraft separation [28], is maintained above 18 meters for the entire transfer duration, despite the fact that no state constraints are applied in the solution procedure.
-The maneuver sequence under the proposed algorithm requires 0.3095 meters per second of total $\Delta V$ expenditure.
-In contrast, [22] reports 0.4934 meters per second of $\Delta V$ expenditure for the same transfer.
-Given that this transfer is performed 20 times over the mission span, this amounts to a total $\Delta V$ saving of 3.68 meters per second, which is $\sim$25% of the total $\Delta V$ budget of the CubeSat.
-The proposed algorithm is additionally capable of incorporating $J_{2}$ and drag perturbations through the linear model formulations in [25], does not require manual tuning of waypoints along the trajectory, and is flexible to changing initial and final conditions.
-The total solver clock time is 0.2675 seconds on a modern computer, which is reasonably achievable in an online manner under the lower computational resources of the VISORS CubeSats.
+Figure 3: VISORS Transfer Trajectory.
+Figure 4: VISORS Transfer Relative Orbital Elements
+The computed VISORS transfer trajectory is displayed in Figure 3, including the RTN trajectory, a subset of the impulse profile, and a subset of the dual norm trajectory.
+Additionally, the relative orbital element trajectory is plotted in Figure 4, along with the RN plane separation, which validates safe spacecraft separation over the transfer duration [8].
+Finally, this scheme is embedded in a high-fidelity simulation, with the inclusion of execution errors, navigation uncertainty, and dynamical effects that are consistent with [11].
+The control dynamical model includes drag and $J_{2}$ effects, as derived in [15], and the transfer is replanned every orbit for the first 9 orbits and every quarter orbit during the final orbit.
+Table I summarizes the performance over 250 Monte Carlo trials.
+Importantly, the proposed approach results in an average of $33.27$ [cm/s] of $\Delta V$ expenditure for the transfer, which substantially improves on the $49.34$ [cm/s] metric reported in [11].
+Given that this transfer is performed 20 times over the mission span, this amounts to an estimated total $\Delta V$ saving of 3.19 meters per second, which is $\sim$25% of the total $\Delta V$ budget of the VISORS CubeSats.
+TABLE I: VISORS Transfer Performance over 250 Monte Carlo Trials with $1\sigma$ Bounds
+Total $\Delta V$ Proposed Solver Time Direct Solver Time
+$33.27\pm 0.72$ [cm/s] $0.1912\pm 0.0480$ [s] $0.4234\pm 0.0263$ [s]
 VI Conclusion
-This paper demonstrated an extension to a class of moment problems that included constraints on the integral of the norm of the input signal over time windows of interest.
-It was shown that the resulting optimal solution remains impulsive and that semi-infinite programming can be applied to global optimality and global convergence.
-This algorithmic framework was applied to challenging spacecraft relative motion control problems, where it was used to derive maneuver sequences that outperformed established techniques in terms of optimality.
-Though this paper has focused on applications to linear time-varying systems, the authors believe that the theoretical framework outlined has significant promise for extending indirect trajectory optimization techniques for impulsively controlled nonlinear systems.
-Future work could also include the incorporation of state constraints and extensions to impulsive covariance control.
+This paper demonstrated an extension to a class of moment problems that includes constraints on the integral of the norm of the input signal over time windows of interest.
+It was shown that the resulting optimal solution remains impulsive and that semi-infinite programming can be applied to obtain optimal solutions with guaranteed convergence.
+This algorithmic framework was applied to challenging spacecraft relative motion control problems, where it was used to compute maneuver sequences that outperformed established techniques in terms of optimality and runtime.